\section*{References}
\begin{thebibliography}{99}
\bibitem{Haake 1990}
 F. Haake, {\it Quantum Signatures of Chaos} 
 (Berlin Heidelberg New York, Springer-Verlag, 1990), 
 Chapters 1, 3, 4, 8 pp 1-11, 33-77, 202-213.
\bibitem{Guhr 1998} 
 T. Guhr T, A. M\"uller-Groeling, and H. Weidenm\"uller,  
 {\it Phys. Rept.} {\bf 299}, 189 (1998). 
\bibitem{Mehta 1990 0}
 M. L. Mehta, {\it Random matrices} 
 (Boston, Academic Press, 1990), Chapters 1, 2, 9 pp 1-54, 182-193.
\bibitem{Ginibre 1965} 
 J. Ginibre, {\it J. Math. Phys.} {\bf 6}, 440 (1965). 
\bibitem{Mehta 1990 1}
 M. L. Mehta, {\it Random matrices} 
 (Boston, Academic Press, 1990), Chapter 15 pp 294-310.
\bibitem{Duras 1996 PRE}
 M. M. Duras, and K. Sokalski,  
 {\it Phys. Rev.} E {\bf 54}, 3142 (1996). 
\bibitem{Duras 1996 thesis}
 M. M. Duras, {\it Finite difference and finite
 element distributions in statistical theory
 of energy levels in quantum systems}
 (PhD thesis, Jagellonian University, Cracow, July 1996). 
\bibitem{Duras 1999 Phys}
 M. M. Duras, and K. Sokalski, 
{\it Physica} {\bf D125}, 260 (1999). 
\bibitem{Duras 1999 Nap}
 M M. Duras, 
 {\it Proceedings of the Sixth International 
 Conference on Squeezed States and Uncertainty Relations,
 24 May-29 May 1999, Naples, Italy} 
 (Greenbelt, Maryland: NASA), at press.
\bibitem{Duras 1996 APPB}
 M. M. Duras, and K. Sokalski, 
 {\it Acta Phys. Pol.} {\bf B27}, 2027 (1996).   
\bibitem{Duras 1997 APPB}
 M. M. Duras, K. Sokalski, and P. Su{\l}kowski, 
{\it Acta Phys. Pol.} {\bf B28} 1023 (1997).
\bibitem{Duras 2000 JOptB}
 M. M. Duras, 
{\it J. Opt. B: Quantum Semiclass. Opt.} {\bf 2}, 287 (2000). 


\end{thebibliography}

\documentstyle[sprocl,epsfig]{article}

\def\be{\begin{equation}}
\def\ee{\end{equation}}
\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}

\begin{document}

\title{COMPLEX-VALUED SECOND DIFFERENCE AS A MEASURE OF
STABILIZATION OF COMPLEX DISSIPATIVE STATISTICAL SYSTEMS: GIRKO ENSEMBLE \\
"Space-time chaos: Characterization, control and synchronization"; 
Proceedings of the International Interdisciplinary School, 
Pamplona, Spain, June 19-23, 2000; 
S. Boccaletti, J. Burguete, W. Gonzalez-Vinas, D. L. Valladares, Eds.; 
World Scientific Publishers, Singapore, 45-52  (2001).
}

\author{MACIEJ M. DURAS$^{\star}$}

\address{$^{\star}$Institute of Physics, Cracow University of Technology, 
ulica Podchor\c{a}\.zych 1, PL-30084 Cracow, Poland
\\E-mail: mduras@riad.usk.pk.edu.pl} 

\maketitle\abstracts{ A quantum statistical system with energy dissipation is studied.
Its statistics is governed by random complex-valued
non-Hermitean Hamiltonians belonging to complex Ginibre ensemble.
The eigenenergies are shown to form stable structure.
Analogy of Wigner and Dyson with system of electrical charges
is drawn.}
\section{Summary}\label{sec-Summary}
A complex quantum system with energy dissipation is considered.
The quantum Hamiltonians $H$ belong the complex Ginibre ensemble.
The complex-valued eigenenergies $Z_{i}$
are random variables.
The second differences $\Delta^{1} Z_{i}$ are also
complex-valued random variables.
The second differences have their
real and imaginary parts and also 
radii (moduli) and main arguments (angles).
For $N$=3 dimensional Ginibre ensemble
the distributions of above random variables are provided
whereas for generic $N$- dimensional Ginibre ensemble 
second difference's, radius's and angle's distributions are
analytically calculated.
The law of homogenization of eigenergies is formulated.
The analogy of Wigner and Dyson
of Coulomb gas of electric charges is studied.
The stabilisation of system of electric charges is dealt with.
\section{Introduction}
\label{sec-introduction}
We study generic quantum statistical systems with energy dissipation.
The quantum Hamiltonian operator $H$ is in given
basis of Hilbert's space a matrix with random
elements $H_{ij}$
\cite{Haake 1990,Guhr 1998,Mehta 1990 0}.
The Hamiltonian $H$ is not Hermitean operator, thus
its eigenenergies $Z_{i}$ are
complex-valued random variables.
We assume that distribution of $H_{ij}$
is governed by Ginibre ensemble
\cite{Haake 1990,Guhr 1998,Ginibre 1965,Mehta 1990 1}.
$H$ belongs to general linear Lie group GL($N$, {\bf C}),
where $N$ is dimension and {\bf C} is complex numbers field.
Since $H$ is not Hermitean, therefore quantum system is
dissipative system. Ginibre ensemble of random matrices
is one of many 
Gaussian Random Matrix ensembles GRME.
The above approach is an example of Random Matrix theory RMT
\cite{Haake 1990,Guhr 1998,Mehta 1990 0}.
The other RMT ensembles are for example
Gaussian orthogonal ensemble GOE, unitary GUE, symplectic GSE,
as well as circular ensembles: orthogonal COE,
unitary CUE, and symplectic CSE.
The distributions of the eigenenergies 
$Z_{1}, ..., Z_{N}$
for $N \times N$ Hamiltonian matrices is given by Jean Ginibre's formula 
\cite{Haake 1990,Guhr 1998,Ginibre 1965,Mehta 1990 1}:
\begin{eqnarray}
& & P(z_{1}, ..., z_{N})=
\label{Ginibre-joint-pdf-eigenvalues} \\
& & =\prod _{j=1}^{N} \frac{1}{\pi \cdot j!} \cdot
\prod _{i<j}^{N} \vert z_{i} - z_{j} \vert^{2} \cdot
\exp (- \sum _{j=1}^{N} \vert z_{j}\vert^{2}),
\nonumber
\end{eqnarray}
where $z_{i}$ are complex-valued sample points
($z_{i} \in {\bf C}$).
For Ginibre ensemble we define complex-valued spacings
$\Delta^{1} Z_{i}$ and second differences $\Delta^{2} Z_{i}$:
\begin{equation}
\Delta^{1} Z_{i}=Z_{i+1}-Z_{i}, i=1, ..., (N-1),
\label{first-diff-def}
\end{equation}
\begin{equation}
\Delta ^{2} Z_{i}=Z_{i+2} - 2Z_{i+1} + Z_{i}, i=1, ..., (N-2).
\label{Ginibre-second-difference-def}
\end{equation}
The $\Delta^{2} Z_{i}$ are extensions of
real-valued second differences
\begin{equation}
\Delta^{2} E_{i}=E_{i+2}-2E_{i+1}+E_{i}, i=1, ..., (N-2),
\label{second-diff-def}
\end{equation}
of adjacent ordered increasingly real-valued energies $E_{i}$
defined for
GOE, GUE, GSE, and Poisson ensemble PE
(where Poisson ensemble is composed of uncorrelated
randomly distributed eigenenergies)
\cite{Duras 1996 PRE,Duras 1996 thesis,Duras 1999 Phys,Duras 1999 Nap,Duras 1996 APPB,Duras 1997 APPB}.

There is an analogy of Coulomb gas of 
unit electric charges pointed out by Eugene Wigner and Freeman Dyson.
A Coulomb gas of $N$ unit charges moving on complex plane (Gauss's plane)
{\bf C} is considered. The vectors of positions
of charges are $z_{i}$ and potential energy of the system is:
\begin{equation}
U(z_{1}, ...,z_{N})=
- \sum_{i<j} \ln \vert z_{i} - z_{j} \vert
+ \frac{1}{2} \sum_{i} \vert z_{i}^{2} \vert. 
\label{Coulomb-potential-energy}
\end{equation}
If gas is in thermodynamical equilibrium at temperature
$T= \frac{1}{2 k_{B}}$ 
($\beta= \frac{1}{k_{B}T}=2$, $k_{B}$ is Boltzmann's constant),
then probability density function of vectors of positions is 
$P(z_{1}, ..., z_{N})$ Eq. (\ref{Ginibre-joint-pdf-eigenvalues}).
Complex eigenenergies $Z_{i}$ of quantum system 
are analogous to vectors of positions of charges of Coulomb gas.
Moreover, complex-valued spacings $\Delta^{1} Z_{i}$
are analogous to vectors of relative positions of electric charges.
Finally, complex-valued
second differences $\Delta^{2} Z_{i}$ 
are analogous to
vectors of relative positions of vectors
of relative positions of electric charges.

The $\Delta ^{2} Z_{i}$ have their real parts
${\rm Re} \Delta ^{2} Z_{i}$,
and imaginary parts
${\rm Im} \Delta ^{2} Z_{i}$, 
as well as radii (moduli)
$\vert \Delta ^{2} Z_{i} \vert$,
and main arguments (angles) ${\rm Arg} \Delta ^{2} Z_{i}$.

\section{Second Difference Distributions}\label{sec-second-difference-pdf}
We define following random variables
for $N$=3 dimensional Ginibre ensemble:
\begin{equation}
Y_{1}=\Delta ^{2} Z_{1}, 
A_{1}= {\rm Re} Y_{1}, B_{1}= {\rm Im} Y_{1},
\label{Ginibre-Y1A1B1-def}
\end{equation}
\begin{equation}
R_{1} = \vert Y_{1} \vert, \Phi_{1}= {\rm Arg} Y_{1},
\label{Ginibre-polar-second-diff-def}
\end{equation}
and for the generic $N$-dimensional Ginibre ensemble
\cite{Duras 2000 JOptB}:
\begin{equation}
W_{1}=\Delta ^{2} Z_{1}, P_{1} = \vert W_{1} \vert, \Psi_{1}= {\rm Arg} W_{1}.
\label{Ginibre-W1P1Psi1-def-N-dim}
\end{equation}
Their distributions for 
are given by following formulae, respectively
\cite{Duras 2000 JOptB}:
\begin{eqnarray}
& & f_{Y_{1}}(y_{1})=f_{(A_{1}, B_{1})}(a_{1}, b_{1})=
\label{Ginibre-marginal-pdf-Y1-def} \\
& & =\frac{1}{576 \pi} [ (a_{1}^{2} + b_{1}^{2})^{2} + 24]
\cdot \exp (- \frac{1}{6} (a_{1}^{2}+b_{1}^{2})),
\nonumber
\end{eqnarray}
is second difference distribution for 3-dimensional Ginibre ensemble.
\begin{equation}
f_{A_{1}}(a_{1})=
\frac{\sqrt{6}}{576 \sqrt{\pi}} (a_{1}^{4}+6a_{1}^{2}+ 51)
\cdot \exp (- \frac{1}{6} a_{1}^{2}),
\label{Ginibre-marginal-pdf-Y1Re-def}
\end{equation}
\begin{equation}
f_{B_{1}}(b_{1})=
\frac{\sqrt{6}}{576 \sqrt{\pi}} (b_{1}^{4}+6b_{1}^{2}+ 51)
\cdot \exp (- \frac{1}{6} b_{1}^{2}),
\label{Ginibre-marginal-pdf-Y1Im-def}
\end{equation}
\begin{eqnarray}
& & f_{R_{1}}(r_{1})=
\label{Ginibre-polar-second-diff-result} \\
& & \Theta(r_{1}) \frac{1}{288}r_{1}(r_{1}^{4}+24) \cdot \exp(- \frac{1}{6} r_{1}^{2}),
\nonumber \\
& & f_{\Phi_{1}}(\phi_{1})= \frac{1}{2 \pi}, \phi_{1} \in [0, 2 \pi],
\nonumber
\end{eqnarray}
are real part's, imaginary part's, modulus's and angle's probability
distributions for 3-dimensional Ginibre ensemble, 
where $\Theta(r_{1})$ is Heaviside (step) function.
Notice, that the angle $\Phi_{1}$ has uniform distribution.
Next, second difference's distribution for $N$-dimensional
Ginibre ensemble reads $P_{3}(w_{1})$ \cite{Duras 2000 JOptB}:
\begin{eqnarray}
& & P_{3}(w_{1})=
\label{W1-pdf-I-result} \\
& & = \pi^{-3} \sum_{j_{1}=0}^{N-1} \sum_{j_{2}=0}^{N-1} \sum_{j_{3}=0}^{N-1}
\frac{1}{j_{1}!j_{2}!j_{3}!}I_{j_{1}j_{2}j_{3}}(w_{1}),
\nonumber \\
& & I_{j_{1}j_{2}j_{3}}(w_{1})=
\label{W1-pdf-I-F} \\
& & = 2^{-2j_{2}} 
\frac{\partial^{j_{1}+j_{2}+j_{3}}}
{\partial^{j_{1}} \lambda_{1} \partial^{j_{2}} \lambda_{2}
\partial^{j_{3}} \lambda_{3}}
F(w_{1},\lambda_{1},\lambda_{2},\lambda_{3}) \vert _{\lambda_{i}=0},
\nonumber
\end{eqnarray}
\begin{eqnarray}
& & F(w_{1},\lambda_{1},\lambda_{2},\lambda_{3})=
\label{W1-pdf-I-F-final} \\
& & = A(\lambda_{1},\lambda_{2},\lambda_{3})
\exp[-B(\lambda_{1},\lambda_{2},\lambda_{3}) \vert w_{1} \vert^{2}],
\nonumber
\end{eqnarray}
\begin{eqnarray}
& & A(\lambda_{1},\lambda_{2},\lambda_{3})=
\label{W1-pdf-I-A} \\
& & =\frac{(2\pi)^{2}}
{(\lambda_{1}+\lambda_{2}-\frac{5}{4}) 
\cdot (\lambda_{1}+\lambda_{3}-\frac{5}{4})-(\lambda_{1}-1)^{2}},
\nonumber \\
& & B(\lambda_{1},\lambda_{2},\lambda_{3})=
\label{W1-pdf-I-B} \\
& & =(\lambda_{1}-1) \cdot \frac{2 \lambda_{1}-\lambda_{2}-\lambda_{3}+\frac{1}{2}}
{2 \lambda_{1}+\lambda_{2}+\lambda_{3}-\frac{9}{2}}.
\nonumber
\end{eqnarray}
Finally,
\begin{eqnarray}
& & f_{P_{1}}(\varrho_{1}) =2 \pi P_{3}(\varrho_{1})=
\label{Ginibre-polar-second-diff-result-N-dim} \\
& & = 2 \pi^{-2} \sum_{j_{1}=0}^{N-1} \sum_{j_{2}=0}^{N-1} \sum_{j_{3}=0}^{N-1}
\frac{1}{j_{1}!j_{2}!j_{3}!}I_{j_{1}j_{2}j_{3}}(\varrho_{1}),
\nonumber \\
& & f_{\Psi_{1}}(\psi_{1})= \frac{1}{2 \pi}, \psi_{1} \in [0, 2 \pi],
\nonumber
\end{eqnarray}
are modulus's and angle's probability
distributions for $N$-dimensional Ginibre ensemble.
Notice, that again the angle $\Psi_{1}$ has uniform distribution.

\section{Conclusions}\label{sect-conclusions}
We compare second difference distributions for different ensembles 
by defining following dimensionless second differences:
\begin{equation}
C_{\beta} = \frac{\Delta^{2} E_{1}}{<S_{\beta}>},
\label{rescaled-second-diff-GOE-GUE-GSE-PE}
\end{equation}
\begin{equation}
X_{1}=\frac{A_{1}}{<R_{1}>},
\label{Ginibre-X1-def} 
\end{equation}
where $<S_{\beta}>$ are
the mean values of spacings 
for GOE(3) ($\beta=1$),
for GUE(3) ($\beta=2$),
for GSE(3) ($\beta=4$), for PE ($\beta=0$)
\cite{Duras 1996 PRE,Duras 1996 thesis,Duras 1999 Phys,Duras 1999 Nap,Duras 1996 APPB,Duras 1997 APPB},
and $<R_{1}>$ is mean value of radius $R_{1}$ 
for $N$=3 dimensional Ginibre ensemble \cite{Duras 2000 JOptB}.

On the basis of comparison of results for
Gaussian ensembles, Poisson ensemble, and Ginibre ensemble
we formulate
\cite{Duras 1996 PRE,Duras 1996 thesis,Duras 1999 Phys,Duras 1999 Nap,Duras 1996 APPB,Duras 1997 APPB,Duras 2000 JOptB}: 
\begin{quote}
{\bf Homogenization Law:} {\it Random eigenenergies of statistical systems
governed by Hamiltonians belonging to Gaussian orthogonal ensemble, 
to Gaussian unitary ensemble, to Gaussian symplectic ensemble,
to Poisson ensemble,
and finally to Ginibre ensemble tend to be homogeneously distributed.}
\end{quote}
It can be restated mathematically as follows:
\begin{quote}
{\it If $H \in$ {\rm GOE, GUE, GSE, PE, Ginibre ensemble}, then
{\rm Prob}($D$)={\rm max}, 
where D is random event corresponding to vanishing of
second difference's probability distributions at the origin.}
\end{quote}
Both of above formulation follow from the fact that
the second differences' distributions assume global maxima at origin
for above ensembles
\cite{Duras 1996 PRE,Duras 1996 thesis,Duras 1999 Phys,Duras 1999 Nap,Duras 1996 APPB,Duras 1997 APPB,Duras 2000 JOptB}.
For Coulomb gas's analogy the vectors of relative positions of vectors
of relative positions of charges statistically most probably vanish.
It means that the vectors of relative positions tend to be equal to each other.
Thus, the relative distances of electric charges are most probably equal.
We call such situation stabilisation
of structure of system of electric charges on complex plane.
 
\section*{Acknowledgements}\label{sect-acknowledgements}
It is my pleasure to most deeply thank Professor Jakub Zakrzewski
for formulating the problem. 
I also thank Professor Antoni Ostoja-Gajewski
for creating optimal environment for scientific work
and Professor W{\l}odzimierz W\'ojcik
for his giving me access to computer facilities.

\end{document}